# A New Model-Independent Precision Test of General Relativity using LISA Bright Standard Sirens

### v1

@@ 8 Conclusion & Discussion
In this study, we introduce a model-independent approach for investigating the frictional term in theories of gravity, capable of identifying deviations from GR in the cosmological scales across any redshift. This analysis is facilitated by bright sirens that will be observable by the upcoming space-based GWs detector LISA, which is designed to operate in the milliHertz frequency range. Building on our previous work [56], we have already demonstrated a model-independent reconstruction of the frictional term using current and forthcoming ground-based GWs detectors, which primarily operate in the hertz to kilohertz frequency spectrum. We propose a novel strategy that integrates three pivotal cosmological measures as functions of redshift: the luminosity distance from GWs sources $D_{l}^{\mathrm{GW}}(z)$, the angular scale of the BAO from galaxy surveys $\theta_{\mathrm{BAO}}(z)$, and the sound horizon $r_{s}$ from CMB observations. This amalgamation provides a unique framework to scrutinize alternate theories of gravity, particularly focusing on the frictional term. Through an integrated analysis combining BAO measurements, sound horizon distances from the CMB, EM counterpart redshifts, and GWs luminosity distances, we enable a robust measurement of cosmological parameters. This approach notably improves our ability to simultaneously determine the Hubble constant ($H_{0}$) and the frictional term as a function of redshift in a model-independent way. We have demonstrated the feasibility of achieving precise measurements of the frictional term with an accuracy ranging from 2.38% to 7.21% up to a redshift of 6, utilizing just a single GW source detected by LISA. Additionally, we achieve a measurement of the Hubble constant with a precision of 1.27%, marginalized over the measurements of deviations in the frictional term, for 4 years of observation time. This assumes that EM are detected for 75% of the events. The accuracy of these measurements scales inversely with the square root of the number of events, which is proportional to the square root of the observation time, expressed as $1/\sqrt{T_{\mathrm{obs}}/\text{years}}$. So, if the mission duration lasts for $10$ years, then about a factor of $2.5$ times decrease in the error bar on $\mathcal{F}(z)$ is feasible. Furthermore, our methodology allows for the reconstruction of the frictional term as a function of redshift without relying on any specific parametric forms, highlighting the strength and flexibility of this model-independent approach in advancing our understanding of cosmic expansion and the underlying forces.
@@
The current planned BAO measurements with detectors such as Vera Rubin Observatoryand Euclid will significantly enhance our understanding by surveying 18,000 square degrees and extending precise measurements within the redshift range of z=1 to z=3. Additionally, DESI focuses on BAO and the evolution of cosmic structures through redshift-space distortions up to redshift $z=3.7$. Despite covering 14,000 square degrees and targeting an accuracy better than 0.5% within the redshift intervals of 0.0 < z < 3.7, there are currently no instruments planned to measure BAO beyond redshift $z=3.7$. This limitation on the BAO measurements impacts our results on the reconstruction of the frictional term as a function of cosmic redshift. This limitation is illustrated in Figures 10 and 11, where we show the results using deeper colors for redshifts up to $z=3.0$. Beyond redshift $z=3.0$, the results are shown in fainter colors to highlight the limitations of our current planned BAO measurements, as we have used LSST-like measurements for our analysis. Improved surveys with deeper reach could potentially overcome these limitations, providing more accurate reconstructions and deeper insights into the frictional term across a wider range of redshifts.
@@
Additionally, our proposed methodology can be adapted to incorporate dark sirens, where accurately determining the redshift is critical. For this, we employ the three-dimensional clustering of GW sources with galaxies to deduce redshifts, as highlighted in prior studies [173, 174, 175, 176, 177]. It is important to note that the current analysis focuses strictly on bright sirens. This focus necessitates the exclusion of certain GW sources such as EMRIs and stellar mass BHBs, which typically do not have EM counterparts. Furthermore, our study limits its scope to redshifts up to $z=8$ due to the increasing difficulty in detecting EM counterparts beyond this point. Incorporating dark sirens, which do not have EM counterparts, could significantly refine the accuracy of our frictional term measurements. The extension of our approach to include dark sirens could lead to more precise determinations of the frictional term, thereby enhancing our understanding of the fundamental principles of gravity.

### v2

@@ 8 Conclusion & Discussion
In this study, we introduce a model-independent approach for investigating the frictional term in theories of gravity, capable of identifying deviations from GR in the cosmological scales across any redshift. This analysis is facilitated by bright sirens that will be observable by the upcoming space-based GWs detector LISA, which is designed to operate in the milliHertz frequency range. Building on our previous work [57, 176], we have already demonstrated a model-independent reconstruction of the frictional term using current and forthcoming ground-based GW detectors, which primarily operate in the hertz to kilohertz frequency spectrum.
@@
We propose a novel strategy that integrates three pivotal cosmological measures as functions of redshift: the luminosity distance from GW sources $\mathrm{D_{L}^{\mathrm{GW}}(z)}$, the angular scale of the BAO from galaxy surveys $\mathrm{\theta_{BAO}(z)}$, and the sound horizon $\mathrm{r_{s}}$ from CMB observations. This amalgamation provides a unique framework to scrutinize alternate theories of gravity, particularly focusing on the frictional term. Through an integrated analysis combining BAO measurements, sound horizon distances, redshiftd from EM counterpart, and GW luminosity distances, we enable a robust measurement of cosmological parameters. This approach notably improves our ability to simultaneously determine the Hubble constant ($\mathrm{H_{0}}$) and the frictional term as a function of redshift in a model-independent way. We have demonstrated the feasibility of achieving precise measurements of the frictional term with an accuracy ranging from $\mathrm{2.38\%}$ to $\mathrm{7.21\%}$ up to a redshift of 6, utilizing just a single GW source detected by LISA. Additionally, we achieve a measurement of the Hubble constant with a precision of $\mathrm{1.27\%}$, marginalized over the measurements of deviations in the frictional term, for 4 years of observation time. This assumes that EM are detected for $\mathrm{75\%}$ of the events. The accuracy of these measurements scales inversely with the square root of the number of events, which is proportional to the square root of the observation time, expressed as $\mathrm{1/\sqrt{T_{\mathrm{obs}}/\text{years}}}$. Furthermore, our methodology allows for the reconstruction of the frictional term as a function of redshift without relying on any specific parametric forms, highlighting the strength and flexibility of this model-independent approach in advancing our understanding of cosmic expansion and the underlying forces.
@@
The current BAO measurements from the Vera Rubin Observatory, Euclid, and DESI, which cover redshifts up to z=3.7. This limitation on the BAO measurements impacts our results on the reconstruction of the frictional term as a function of cosmic redshift. This limitation is illustrated in Figures 11 and 12, where we show the results using deeper colors for redshifts up to $\mathrm{z=3.0}$. Beyond redshift $\mathrm{z=3.0}$, the results are shown in fainter colors to highlight the limitations of our current planned BAO measurements, as we have used LSST-like measurements for our analysis. Improved surveys with deeper reach could potentially overcome these limitations, providing more accurate reconstructions and deeper insights into the frictional term across a wider range of redshifts.
@@
Additionally, our proposed methodology can be adapted to incorporate dark sirens, where accurately determining the redshift is critical. For this, the three-dimensional clustering of GW sources with galaxies is employed to deduce redshifts, as highlighted in prior studies. [177, 178, 179, 180, 181, 176]. It is important to note that the current analysis focuses strictly on bright sirens. This focus necessitates the exclusion of certain GW sources such as EMRIs and stellar mass BHBs, which typically do not have EM counterparts. Furthermore, our study limits its scope to redshifts up to $\mathrm{z=6}$ due to the increasing difficulty in detecting EM counterparts beyond this point. Incorporating dark sirens, which do not have EM counterparts, could significantly refine the accuracy of our frictional term measurements. The extension of our approach to include dark sirens could lead to more precise determinations of the frictional term, thereby enhancing our understanding of the fundamental principles of gravity.
@@ Acknowledgements
Appendix A Impact of Electromagnetic Counterpart Detection on $\mathrm{\mathcal{F}(z)}$ and $\mathrm{H_{0}}$ Measurements
In this section, we present a few cases of bright sirens detectable from LISA and summarize the best measurements of $\mathrm{\mathcal{F}(z)}$ and the combined error on $\mathrm{H_{0}}$ for three different models: Light Seed, Heavy Seed, and Heavy Seed + Delay. Each table provides results for various event selections, including 75%, 50%, 25%, and 10% of selected events. The tables detail the maximum redshift up to which a source is detected in LISA ($\mathrm{z_{max}}$), the total number of GW events ($\mathrm{N_{GW}}$) up to $\mathrm{z_{max}}$, and the associated precision on $\mathrm{\mathcal{F}(z)}$ and $H_{0}$. For $\mathrm{\mathcal{F}(z)}$ we quote the minimum uncertainty and corresponding redshift, and for the Hubble constant, we mention the uncertainty after combining $\mathrm{N_{GW}}$ events.
Tables 3, 4, and 5 present the results for the Light Seed model, Heavy Seed model and the Heavy Seed + Delay model respectively. For each model, the tables highlight the least percentage error in $\mathrm{\mathcal{F}(z)}$ at specific redshifts, and the overall combined error on $\mathrm{H_{0}}$. Note that for the 10% selection case, where no events are available, it is denoted as "–". These tables illustrate how the precision of $\mathrm{\mathcal{F}(z)}$ and $\mathrm{H_{0}}$ varies with the number of selected events and their redshift range, reflecting the impact of different selection thresholds on measurement accuracy.
Bright events $\mathrm{z_{max}}$ $\mathrm{N_{GW}}$ Least uncertainty on $\mathrm{\mathcal{F}(z)}$ Combined uncertainty on $\mathrm{H_{0}}$
75% Selected 6 73 2.65% at z=1.2 1.62%
50% Selected 6 49 2.65% at z=1.2 2.05%
25% Selected 6 23 3.57% at z=2.0 3.40%
10% Selected 5.2 6 5.02% at z=3.2 6.42%
Table 3: Table showing the best measurement of $\mathrm{\mathcal{F}(z)}$ and the combined error on $\mathrm{H_{0}}$ for the Light Seed model. The table summarizes results for various event selections, detailing the maximum redshift ($\mathrm{z_{max}}$), the number of GW events ($\mathrm{N_{GW}}$), and the associated precision for 75%, 50%, 25%, and 10% of selected events. For each selection, the table highlights the least percentage error in $\mathrm{\mathcal{F}(z)}$ at a specific redshift and the overall combined error on $\mathrm{H_{0}}$. These results illustrate how the precision of $\mathrm{\mathcal{F}(z)}$ and $\mathrm{H_{0}}$ varies with the number of selected events and their redshift range, reflecting the impact of different selection thresholds on measurement accuracy.
Bright events $\mathrm{z_{max}}$ $\mathrm{N_{GW}}$ Least uncertainty on $\mathrm{\mathcal{F}(z)}$ Combined uncertainty on $\mathrm{H_{0}}$
75% Selected 6 98 2.62% at z=1.2 1.27%
50% Selected 6 66 2.62% at z=1.2 1.56%
25% Selected 6 32 3.02% at z=1.6 2.60%
10% Selected 6 9 4.51% at z=2.8 4.50%
Table 4: Table showing the best measurement of $\mathrm{\mathcal{F}(z)}$ and the combined error on $\mathrm{H_{0}}$ for the Heavy Seed model. The table summarizes results for various event selections, detailing the maximum redshift ($\mathrm{z_{max}}$), the number of GW events ($\mathrm{N_{GW}}$), and the associated precision for 75%, 50%, 25%, and 10% of selected events. For each selection, the table highlights the least percentage error in $\mathrm{\mathcal{F}(z)}$ at a specific redshift and the overall combined error on $\mathrm{H_{0}}$. These results illustrate how the precision of $\mathrm{\mathcal{F}(z)}$ and $\mathrm{H_{0}}$ varies with the number of selected events and their redshift range, reflecting the impact of different selection thresholds on measurement accuracy.
Bright events $\mathrm{z_{max}}$ $\mathrm{N_{GW}}$ Least uncertainty on $\mathrm{\mathcal{F}(z)}$ Combined uncertainty on $\mathrm{H_{0}}$
75% Selected 6 31 2.63% at z=1.2 2.42%
50% Selected 6 21 2.63% at z=1.2 3.20%
25% Selected 5.6 9 4.26% at z=2.4 6.82%
10% Selected
Table 5: Table showing the best measurement of $\mathrm{\mathcal{F}(z)}$ and the combined error on $\mathrm{H_{0}}$ for the Heavy Seed + Delay model. The table summarizes results for various event selections, detailing the maximum redshift ($\mathrm{z_{max}}$), the number of GW events ($\mathrm{N_{GW}}$), and the associated precision for 75%, 50%, 25%, and 10% of selected events. For each selection, the table highlights the least percentage error in $\mathrm{\mathcal{F}(z)}$ at a specific redshift and the overall combined error on $\mathrm{H_{0}}$. Note that for the 10% case, there are no events, so it is denoted as "–". These results illustrate how the precision of $\mathrm{\mathcal{F}(z)}$ and $\mathrm{H_{0}}$ varies with the number of selected events and their redshift range, reflecting the impact of different selection thresholds on measurement accuracy.
Appendix B Measurements for Different Mass Ratio Distributions
In this section, we extend our analysis of $\mathcal{F}(z)$ and $H_{0}$ using LISA bright sirens by considering different mass ratio distributions for binary mergers. Previously, we presented results assuming mass ratio $q=1$. Here, we explore the impact of varying mass ratios on the measurements of $\mathcal{F}(z)$ and $H_{0}$. Specifically, we investigate mass ratio distributions following $\mathrm{q^{2}}$, where $\mathrm{q}$ is defined as the ratio of the component masses and ranges from $\mathrm{10^{-3}}$ to 1. We consider two selection criteria for these distributions: 75% and 50%, corresponding to the different proportions of systems retained based on their mass ratios. This approach allows us to understand how variations in the mass ratio distribution affect the detectability and measurement precision of bright sirens.
In Table 6 we provides a summary of the total number of events and the number of detected events for different merger models with an SNR threshold of 50, categorized by the mass ratio selection criteria of 75% and 50%. The results show that the distribution of mass ratios does not significantly affect the number of detected events, as our study is limited to sufficiently low redshifts where most events are very bright. Comparing these findings with the results in Table 2, we find that the number of detected events differs by approximately 2-7%, depending on the model. This variation is most pronounced in the high-redshift portion of our analysis.
Model $T_{obs}$ (years) Total Events Selected (75%) Selected (50%) 1 Event Detectability
LS 4 158 68 44 6.82%
HS 4 147 96 64 4.69%
HSD 10 49 29 20 15.01%
Table 6: Number of total and detected events for different merger models Light Seed (LS), Heavy Seed (HS), and Heavy Seed+Delay (HSD), with a threshold SNR of 50. The table includes results based on 75% and 50% selection criteria, as well as the percentage of detectability of single events at each integer redshift from 1 to 6 (1 Event Detectability).
Figures 15 and 16 illustrate the results on the feasibility of inferring the parameter $\mathcal{F}(z)$ using the $\mathrm{q^{2}}$ mass ratio distribution with LISA. Figure 15 presents the precision in reconstructing the redshift evolution of the Planck mass, $\mathcal{F}(z)$, highlighting LISA’s measurement capabilities up to a redshift of 6, with reduced accuracy beyond a redshift of 3.0 due to limitations in BAO scale measurements. Figure 16 shows the posterior distribution of the Hubble constant ($\mathrm{H_{0}}$) for events up to a redshift of $z=6.0$, assuming that EM counterparts are detected for 75% of the events.
Here, we have used a power-law distribution proportional to $\mathrm{q}^{\alpha}$ with $\alpha=2$. The measurements of $\mathrm{F(z)}$ and $\mathrm{H_{0}}$ will improve for any $\alpha>2$, as the area under the curve of these distributions increases when $\mathrm{q}$ approaches 1. This means that the probability of obtaining nearly equal mass binaries is higher, leading to a higher SNR and, consequently, more events. Conversely, for $\alpha<2$, the measurements become less accurate because the area under the curve decreases, reducing the likelihood of nearly equal mass binaries, resulting in a lower SNR and fewer events.
Figure 15: This plot illustrates the precision in reconstructing the redshift evolution of the Planck mass, represented by $\mathrm{\mathcal{F}(z)}$, for the Light Seed, Heavy Seed, and Heavy Seed+delay models using the upcoming space-based LISA observatory. The observation time is 4 years for the Light Seed and Heavy Seed models and 10 years for the Heavy Seed+delay model, assuming the $\mathrm{q^{2}}$ mass ratio distribution and that EM counterparts are detected for 75% of the events. The plot highlights LISA’s accuracy in measuring $\mathrm{\mathcal{F}(z)}$ up to a redshift of 6. Beyond a redshift of 3.0, a fainter color is used to indicate that the BAO scale cannot be measured beyond this redshift with the currently planned galaxy surveys.
[…]